OpenStudy (anonymous):

Define x and y implicitly as differential functions x=f(t), y=g(t), find the slope of the curve x=f(t), y=g(t) at the given value of t x=√5-(√t), y(t-1)=lny, t=1

OpenStudy (anonymous):

It's just finding derivatives ^.^ I trust you can find \[\large \frac{dx}{dt}\] without difficulty?

OpenStudy (anonymous):

yikes... trickier than I thought :( Maybe you can introduce a new parameter, t-1 instead?

OpenStudy (abb0t):

Can you re-write the function using LaTex? I am kind of unsure of what it is..

OpenStudy (abb0t):

You need to take the derivative of the function with respect to, \(t\) for both \(x\) and \(y\). That means \(\frac{dx}{dt}\) and \(\frac{dy}{dt}\)...your slope is therefore \(\huge \frac{\frac{dy}{dt}}{\frac{dx}{dt}}\)

OpenStudy (abb0t):

evaluate the slope: \(\huge \frac{\frac{dy}{dt}}{\frac{dx}{dt}}\) at t=1

OpenStudy (anonymous):

I think I panicked ^.^